I have 3 cups with 3 marbles each; one black and two white. If I draw a marble from each cup, what is the probability that I will draw AT LEAST ONE white marble? With one cup it's 66%. If I have 100 cups, the probability that I will draw at least 1 white marble is nearly 190%. How do you calculate the actual probability of drawing at least one white marble, given N cups?
1
Expert's answer
2019-11-13T11:19:19-0500

Let "B" be the event "draw a black marble from a cup". Then


"1 \ cup:P(B)={1 \over 1+2}={1 \over 3}"

"2 \ cups:P(BB)={1 \over 3}\cdot{1 \over 3}=({1 \over 3})^2"

"3 \ cups:P(BBB)=({1 \over 3})^3"

"..."

"N \ cups: P(B\ \text{n times})=({1 \over 3})^N"

"P(\text{at least 1 white marble if I have N cups)}="

"=1-P(B\ \text{n times})=1-({1 \over 3})^N"

"100\ cups:P(\text{at least 1 white marble if I have 100 cups)}="

"=1-({1 \over 3})^{100}\approx"

"\approx0.99999999999999999999999999999999999999999999999806"